OpenStudy (anonymous):

what is the slope of the line that passes through the points (4,1) and (1,2)

OpenStudy (anonymous):

Use the gradient formula to find out the gradient first. Then use the gradient and one point and substitute it back into the general linear equation y= mx + c where m is your gradient and c is your y-intercept.

OpenStudy (anonymous):

okay now what?

OpenStudy (anonymous):

what do u mean? what step are u up to?

OpenStudy (anonymous):

im doing the equation that chelseachiodos told me

OpenStudy (anonymous):

I will repost it if you'd like

OpenStudy (anonymous):

@ChelseaChiodos, you can guide @blah500 then.

OpenStudy (anonymous):

ok, you have (x1,y1), (x2,y2) answer is Y-y1=m(X-x1) where m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Alright. You get the slope using the slope formula from the 2 equations provided like I said before.Then you take the slope that you got and plug it into the point-slope formula which is y-y1=m(x-x1).

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i'm sorry i explain line formula, for slop you just compute m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

thats okay thanks though

OpenStudy (anonymous):

do you want to help you?